OpenStudy (anonymous):

proof integration sec^2 =tan x

OpenStudy (anonymous):

You mean prove: \[ \int \sec^2x \;dx=\tan x+C \]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Are we allowed to use the Fundy Theo Calculus?

OpenStudy (anonymous):

what is that??? first time i heard of it

OpenStudy (anonymous):

Fundamental Theorem of Calculus states that an indefinite integral is just the anti-derivative

OpenStudy (anonymous):

What are we allowed to use to prove it?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Okay then, differentiate \(\tan(x)\) and you win!

OpenStudy (anonymous):

if from differentiate from tan x i have done it... but my lecture told me he want from integratretion of sec^2

OpenStudy (anonymous):

Hmmm, tough!

OpenStudy (anonymous):

\(\sec^2 x = 1/\sin^2 x\)

OpenStudy (anonymous):

\[\sec ^{2}x= \frac{ 1 }{ \cos ^{2} }\]

OpenStudy (anonymous):

Right... I messed up lol.

OpenStudy (anonymous):

Wait....

OpenStudy (anonymous):

If we can't do derivative of \(\tan\) then we cant do derivative of any other function right?

OpenStudy (anonymous):

yes.. you right. you only can use integration method to proof it

OpenStudy (anonymous):

Do we have to use the definition of integral? That is too tough!

OpenStudy (anonymous):

yes ..its very taugh... my lecturer said no pain no gain

OpenStudy (anonymous):

I can't really help you because I don't actually know what we are allowed to use...

OpenStudy (anonymous):

The definition of the integral is: \[ \int\limits_a^b f(x) dx \equiv \lim_{\Delta x\to 0}\sum_if(x_i^*)\Delta x_i \]Where the interval \([a,b]\) is partitioned into intervals, \(x_i^*\) is any element in the \(i\)th partition, \(\Delta x_i\) is the length of the \(i\)th partition, and \(\Delta x\) is the largest \(\Delta x_i\) (the largest partition length). This definition can become less generalized by coming up with a method partitioning and choosing \(x^*\). Consider: - Let \(n\) be the number of partitions - Give each partition have equal length: \(\Delta x = \frac{b-a}{n}\) - Choose the right most element in the partition: \(x_i^*=i\Delta x\). Notice that as \(n\to \infty,\;\Delta x \to 0\). Our less generalized definition is: \[ \int\limits_a^b f(x) dx \equiv \lim_{n\to \infty}\sum_i^nf\left(\frac{(b-a)i}{n}\right)\frac{b-a}{n} \]

hartnn (hartnn):

i can think of this, put t= tan x dt = sec^2 x dx \(\int \sec^2xdx = \int dt = t+c = \tan x+c\) i have basically used differentiation but in the disguise of substitution. maybe it'll work :P